Two metal rods in a factory are oppositely charged and placed 8.9 cm apart. One rod has a charge of +7.5 × 10-7 C and the other has a charge of –5.1 × 10-5 C. What is the force between the rods? Is it an attractive or repulsive force?
1
Expert's answer
2020-12-06T17:31:35-0500

Force is Attractive.

Since lengths are not specified, let's consider them to be equivalent to point charges.

Applying,

"F= \frac{k\times Q_{1} \times Q_{2}}{R^{2}}"

On substitution and solving,

F = 43.46 N